# FPT approximations for Capacitated Sum of Radii and Diameters

### v1

@@ 5.2 Hardness of FPT-approximation for Uniform CapSoR
Let $\phi$ be an instance of Max3333-SAT5555 on $N$ variables and $M$ clauses. Let $\varepsilon^{\prime}>0$ be a constant and $k\leq N$ be a positive integer. Furthermore, let $w=\max\{10N,2^{\lceil N/k\rceil}\}$. Then, in ${\rm poly}(k2^{N/k})$ time one can compute an instance $\mathcal{I}=((P,\delta),k,U)$ of uniform CapSoR with $|P|=k(2^{\lceil N/k\rceil}+w)+M$ points and uniform capacity constrain $U=w+2^{\lceil N/k\rceil}+5\lceil N/k\rceil$, such that
  • If $\phi$ is satisfiable then there exist a feasible solution $(S,\sigma)$ with ${\rm cost}(S,\sigma)\leq k$
  • If every assignment satisfies less than $(1-\varepsilon^{\prime})M$ clauses, then for every feasible solution $(S,\sigma)$ it holds that ${\rm cost}(S,\sigma)>(1+\varepsilon^{\prime}/66)k$.
@@
Given a Max3333-SAT5555 instance $\phi$ on $N$ variables and $M$ clauses and a positive integer $k\leq n$, partition the variables (arbitrarily) into $k$ groups $G_{1},\cdots,G_{k}$ such that each $G_{i}$ has at most $\lceil N/k\rceil$ variables. We create the point set $P=P_{1}\cup P_{2}\cup P_{3}$ of CapSoR as follows.
  • Assignment points of $G_{i}$. For every group $G_{i}$ and every assignment $a(i,j)$ to the variables of $G_{i}$, we create a point $p(i,j)$ in $P$. Let $P^{i}_{1}$ denote the set of all assignment points of $G_{i}$. If $|G_{i}|<2^{\lceil N/k\rceil}$, add dummy points to $P^{i}_{1}$ to make $|P^{i}_{1}|=2^{\lceil N/k\rceil}$. Let $P_{1}=\cup_{i\in[k]}P^{i}_{1}$.
  • Clause points. We create a point in $P$ for every clause in $\phi$. Let $P_{2}$ denote the set of all clause points.
  • Auxiliary points. We create $w=\max\{10N,2^{\lceil N/k\rceil}\}$ auxiliary points for every $G_{i}$. Let $P^{i}_{3}$ denote the set of all auxiliary points of $G_{i}$ and let $P_{3}=\cup_{i\in[k]}P^{i}_{3}$.
Note that $|P|=k(2^{\lceil N/k\rceil}+w)+M$.
Distance function $\delta$. It is convenient to think of the distance function $\delta$ as the shortest path metric of an unweighted graph $G=(V,E)$, where $V=P$, and add edges in $E$ between $(1)$ every pair of clause points, $(2)$ every pair of vertices in $P^{i}_{1}\cup P^{i}_{3}$ for every $i$, $(3)$ between every assignment point and and every clause point it satisfies. Note that the distance aspect ratio $\Delta$ of $(P,\delta)$ is $5$. See Figure 4 for an illustration.
Uniform capacity. Set $U=w+2^{\lceil N/k\rceil}+5\lceil N/k\rceil$.
Figure 4: A pictorial illustration of the graph $G$ obtained from the reduction, showing assignment points, auxiliary points, and clause points, along with partial edges.
@@
If $\phi$ is satisfiable, then there exists a feasible solution $(S,\sigma)$ to $\mathcal{I}$ such that ${\rm cost}(S,\sigma)\leq k$.
@@
Suppose $\alpha:X\rightarrow\{0,1\}$ is a satisfying assignment for $\phi$. For each $i\in[k]$, pick $s_{i}\in P^{i}_{1}$ such that $s_{i}$ corresponds to the assignment to the variables of $G_{i}$ given by $\alpha$. Let $S=\{s_{i}\}_{i\in[k]}$ and note that $|S|=k$. Now, we define the clusters $\sigma:P\rightarrow S$ as follows. For $p\in P^{i}_{1}\cup P^{i}_{3}$, set $\sigma(p)=s_{i}$. For every clause point $p\in P_{2}$, set $\sigma(p)=s_{i}$ (breaking ties for $s_{i}$ arbitrary) such that the assignment corresponding to $s_{i}$ satisfies the clause corresponding to $p$.
First note that, for $s_{i}\in S$, we have $|\sigma^{-1}(s_{i})|\leq 2^{\lceil N/k\rceil}+w+5\lceil N/k\rceil=U$ since each $s_{i}$ is assigned at most $5\lceil N/k\rceil$ clause points corresponding to the clauses satisfied the assignment of $s_{i}$. Finally, note that ${\rm rad}(s_{i})=1$ since the points in $\sigma^{-1}(s_{i})$ are neighbors of $s_{i}$ (at distance $1$ from $s_{i}$). Hence, ${\rm cost}(S,\sigma)=k$. ∎
Lemma 15.
Suppose that every assignment satisfies less than $(1-\varepsilon^{\prime})M$ clauses in $\phi$. Then for every feasible solution $(S,\sigma)$, it holds that ${\rm cost}(S,\sigma)>(1+\varepsilon)k$, for $\varepsilon=\varepsilon^{\prime}/66$.
Proof.
@@
We prove the contraposition. Suppose there exists $S\subseteq P,|S|\leq k$ and a feasible assignment $\sigma:P\rightarrow S$ such that ${\rm cost}(S,\sigma)\leq(1+\varepsilon)k$. We will show an assignment that satisfies at least $(1-\varepsilon^{\prime})M$ clauses in $\phi$. To this end, first, we show that $|S|=k$ and there are no zero radius centers in $S$.
Claim 16.
$|S|=k$ and for every $s\in S$, we have that ${\rm rad}(s)\geq 1$.
Proof.
First, suppose $|S|\leq(k-1)$. Then, the number of points served by $S$ is
$|\sigma^{-1}(S)|\leq(k-1)U=(k-1)(2^{\lceil N/k\rceil}+w+5\lceil N/k\rceil)<k(2^{\lceil N/k\rceil}+w)+5k\lceil N/k\rceil-w<|P|~{},$
since $5k\lceil N/k\rceil\leq 10N\leq w$. Now, assume that there is a zero radius center in $S$. Hence, the number of points served by $S$ is
$\leq(k-1)U+1=(k-1)(2^{\lceil N/k\rceil}+w+5\lceil N/k\rceil)+1<k(2^{\lceil N/k\rceil}+w)+5k\lceil N/k\rceil-w<|P|,$
since $\lceil N/k\rceil\geq 1$. ∎
We will need the following definitions.
Definition 14 (Block).
For $i\in[k]$, the $i^{th}$ block $B_{i}$ is the set of points in $P^{i}_{1}\cup P^{i}_{3}$. Block $B_{i}$ is said to be covered by $S$ if $|B_{i}\cap S|\geq 1$, otherwise it is uncovered by $S$. Block $B_{i}$ is a cheating block with respect to $S$ if $|B_{i}\cap S|\geq 2$. Accordingly, $B_{i}$ is a good block if $B_{i}$ is covered by $S$ and is non-cheating, i.e., $|B_{i}\cap S|=1$. We say a center $s\in S$ cheating if $s$ belongs to a cheating block. $s\in S$ is a bad center if $s\in P_{2}$. Finally, a center is good if it is neither cheating nor bad.
Without loss of generality, we assume that for every good block $B_{i}$, we have $|S\cap P^{i}_{1}|=1$, i.e., $S$ contains a center only from the assignment points of $B_{i}$. To see this, suppose $S$ contains a center $s\in P^{i}_{3}$. Then, note that since $P^{i}_{1}$ separates $s$ from the rest of the graph and the neighbors of $s$ are subset of the neighbors of any point $s^{\prime}\in P^{i}_{1}$, $s^{\prime}$ can serve the points served by $s$ within the radius of $s$. Hence we can replace $s$ with $s^{\prime}$ in $S$ without increasing its cost and violating the capacities.
Definitions:
$\boldsymbol{\ell_{0}}$
: number of uncovered blocks.
$\boldsymbol{\ell_{1}}$: number of of cheating blocks.
$\boldsymbol{\ell_{2}}$: number of cheating centers.
$\boldsymbol{\ell_{3}}$: number of bad centers.
$\boldsymbol{\ell_{4}}$: number of good centers in $S$.
Observations:
$\ell_{2}\geq\ell_{1}$
$\ell_{0}=\ell_{2}-\ell_{1}+\ell_{3}$
$k=|S|=\ell_{2}+\ell_{3}+\ell_{4}$
The next claim says that the number of uncovered blocks by $S$ are bounded.
Claim 17.
$\ell_{0}\leq 2\varepsilon k$
Proof.
Suppose the number of uncovered blocks $\ell_{0}>2\varepsilon k$. Consider an0 uncovered block $B_{i},i\in[k]$. Then, note that any center in $S$ that serves points of $B_{i}$ must have radius at least $2$. Since $U$ is the uniform capacity, there are at least $t=\ell_{0}(2^{\lceil N/k\rceil}+w)/U$ centers in $S$ that have radius at least $2$. Hence, using 16, we have
${\rm cost}(S,\sigma)$ $\geq 2t+(k-t)$
$=k+\frac{\ell_{0}(2^{\lceil N/k\rceil}+w)}{U}$
$\geq k+\ell_{0}/2\qquad\text{ since }U\leq 2(2^{\lceil N/k\rceil}+w)$
$>(1+\varepsilon)k,$
where in the second last inequality we used that $5{\lceil N/k\rceil}\leq w$ and in the last inequality we used $\ell_{0}>2\varepsilon k$. ∎
Now, we will massage $S$ to get $S^{\prime}$ such that every block is good with respect to $S^{\prime}$ and the cost of $S^{\prime}$ is not much larger than that of $S$.
Claim 18.
Given a feasible solution $(S,\sigma)$ with ${\rm cost}(S,\sigma)\leq(1+\varepsilon)k$, we can obtain another feasible solution $(S^{\prime},\sigma^{\prime})$ such that ${\rm cost}(S^{\prime},\sigma^{\prime})\leq(1+11\varepsilon)k$ and every block is good with respect to $S^{\prime}$.
Proof.
The idea is simple - redistribute the cheating and bad centers in $S$ so that every block is good. Since, the number of such centers is small and the diameter of $(P,\delta)$ is bounded, the new solution $S^{\prime}$ also has bounded cost. More formally, for every $s\in S$, we add $\gamma(s)\in P$ to $S^{\prime}$ as follows. For every good block $B_{i}$, we add the center $S\cap B_{i}$ to $S^{\prime}$. Hence, we add $\ell_{4}$ centers to $S^{\prime}$ and thus, $S^{\prime}$ covers $\ell_{4}$ blocks. For every cheating block $B_{i}$ with respect to $S$, we arbitrarily select a center in $s_{i}\in S\cap B_{i}$ as the leader for $B_{i}$ and call the other centers in $S\cap B_{i}$ as the excess centers in $B_{i}$. We add all the leaders from $S$ to $S^{\prime}$. Hence, we add $\ell_{1}$ many points to $S^{\prime}$ and thus $S^{\prime}$ covers $\ell_{1}+\ell_{4}$ blocks. Now, we have to add $k-(\ell_{1}+\ell_{4})$ centers to $S^{\prime}$ and we have to cover $k-(\ell_{1}+\ell_{4})$ blocks. For every bad center $s\in S$, we select an arbitrary uncovered block $B_{j}$ and add an arbitrary point $\gamma(s)\in P^{j}_{1}$ to $S^{\prime}$. Thus, we add $\ell_{3}$ many points to $S^{\prime}$ and hence $S^{\prime}$ covers $\ell_{3}$ more blocks. For every excess center $s\in S$, we select an arbitrary uncovered block $B_{j}$ and add an arbitrary point $\gamma(s)\in P^{j}_{1}$ to $S^{\prime}$. Thus, we add $\ell_{2}-\ell_{1}$ many points to $S^{\prime}$ and hence $S^{\prime}$ covers $\ell_{2}-\ell_{1}$ more blocks. Now, $|S^{\prime}|=\ell_{1}+\ell_{4}+\ell_{3}+(\ell_{2}-\ell_{1})=\ell_{2}+\ell_{3}+\ell_{4}=k$, which is also the number of covered blocks.
Next, we define $\sigma^{\prime}$ as follows. For every $\gamma(s)\in S^{\prime}$, we set $\sigma^{\prime}(p)=\gamma(s)$ for every $p\in\sigma^{-1}(s)$. Thus, $(S^{\prime},\sigma^{\prime})$ is a feasible solution. Furthermore, since we redistributed $\ell_{2}-\ell_{1}+\ell_{3}$ centers in $S$ to obtain $S^{\prime}$, we have,
${\rm cost}(S^{\prime},\sigma^{\prime})$ $\leq{\rm cost}(S,\sigma)+5\cdot(\ell_{2}-\ell_{1}+\ell_{3})$
$\leq(1+\varepsilon)k+10\varepsilon k$
$\leq(1+11\varepsilon)k,$
using 17 and the facts: $\Delta=5$, $\ell_{0}=\ell_{2}-\ell_{1}+\ell_{3}$. ∎
Now, let $S^{\prime}_{g}$ denote the set of centers in $S^{\prime}$ with radius $1$ and let $S^{\prime}_{b}=S^{\prime}\setminus S_{g}$. Let $B_{g}$ and $B_{b}$ be the set of blocks containing centers in $S^{\prime}_{g}$ and $S^{\prime}_{b}$, respectively. Next claim says that $S^{\prime}_{b}$ serves small number of clause points.
Claim 19.
Let $\tilde{M}$ be the number of clause points served by the centers in $S^{\prime}_{b}$. Then, $\tilde{M}\leq 66\varepsilon M$.
Proof.
First note that $|S^{\prime}_{b}|\leq 11\varepsilon k$ otherwise ${\rm cost}(S^{\prime})\geq 2|S^{\prime}_{b}|+(k-|S^{\prime}_{b}|)=k+|S^{\prime}_{b}|>(1+11\varepsilon)k$. As the centers in $S^{\prime}_{g}$ have radius $1$, they can not serve points from blocks in $B_{b}$. Thus, the centers in $S^{\prime}_{b}$ serve at least $|S^{\prime}_{b}|(2^{\lceil N/k\rceil}+w)+\tilde{M}$ points. Thus, there is a center in $s^{\prime}\in S^{\prime}_{b}$ that serves at least $(2^{\lceil N/k\rceil}+w)+M^{\prime}/|S^{\prime}_{b}|$ points. Hence, using $M=5N/3$, we have,
$w+2^{\lceil N/k\rceil}+5\lceil N/k\rceil=U\geq|\sigma^{\prime-1}(s^{\prime})|\geq(2^{\lceil N/k\rceil}+w)+\frac{\tilde{M}}{|S^{\prime}_{b}|}$
it follows that $\tilde{M}\leq|S^{\prime}_{b}|\cdot 5\lceil\frac{n}{k}\rceil<11\varepsilon k\cdot(5\cdot\frac{N}{k}+5)\leq 110\varepsilon N=66\varepsilon M$. ∎
Finally, we are ready to show an assignment $\alpha_{S^{\prime}}:X\rightarrow\{0,1\}$ that satisfies at least $(1-\varepsilon^{\prime})M$ clauses in $\phi$. Let $S^{\prime\prime}_{g}\subseteq S^{\prime}_{g}$ consists of centers in $S^{\prime}_{g}$ that correspond to valid assignment points. Consider $s_{i}\in S^{\prime\prime}_{g}$ corresponding to the assignment $\beta_{1}\cdots\beta_{\lceil N/k\rceil}$ to the variables $x^{i}_{1},\cdots,x^{i}_{\lceil N/k\rceil}$ of $G_{i}$. Then, define $\alpha_{S^{\prime}}(x^{i}_{j})=\beta_{j}$ for $j\in{\lceil N/k\rceil}$. Define $\alpha_{S^{\prime}}$ arbitrary for other variables. Then, since centers in $S^{\prime\prime}_{g}$ serve at least $(1-66\varepsilon)M$ clause points at distance $1$, we have that $\alpha_{S^{\prime}}$ satisfy at least $(1-66\varepsilon)M=(1-\varepsilon^{\prime})M$ clauses in $\phi$. ∎
Proof of Theorem 2.
Let $\phi$ be the instance Max3333-SAT5555 on $N$ variables and $M$ clauses obtained from Theorem 8 and let $\varepsilon^{\prime}$ be the corresponding parameter. Suppose there is algorithm $\mathcal{A}$ that takes an instance $\mathcal{I}=((P,\delta),k,U)$ of Uniform CapSoR as input and finds a $(1+\varepsilon^{\prime}/66)$-approximate solution, in time $f(k)n^{k/h(k)}$, for some non-decreasing and unbounded function $h$. We will use $\mathcal{A}$ to decide whether $\phi$ has a satisfying assignment or all assignments satisfy less than $(1-\varepsilon^{\prime})M$ clauses in time $2^{o(n)}$, refuting gap-ETH. Without loss of generality, we assume that $f$ is non-decreasing and unbounded, and that $f(k)\geq 2^{k}$. Towards this, we will use reduce $\phi$ to an instance $\mathcal{I}=((P,\delta),k,U)$ of Uniform CapSoR using Lemma 13 for a particularly chosen value of $k$ as follows. We choose $k$ to be the largest value such that $f(k)\leq N$ and computing such $k$ can be done in $O(N)$ time [CFK+15]. Thus, the value of $k$ selected depends on $N$, and hence let $k=g(N)$ for some computable function $g$ that is non-decreasing and unbounded. Furthermore, using the facts $f(k)\geq 2^{k}\geq k$ and $2^{k}\leq f(g(N))\leq N$, we have that $g(N)\leq N$ and $k\leq 2\log N$. Let $\mathcal{I}=((P,\delta),k,U)$ be the instance of Uniform CapSoR obtained using Lemma 13 on $\phi$ and $k=g(N)$. We now run Algorithm $\mathcal{A}$ on $\mathcal{I}$ to decide whether $\phi$ has a satisfying assignment or all assignments satisfy less than $(1-\varepsilon^{\prime})M$ clauses. The total running time of our algorithm is bounded by
$(k2^{N/k}+M)^{O(1)}+f(k)|P|^{o(k)}$ $\leq(k2^{N/k}+M)^{O(1)}+f(k)(k(2^{2N/k}+w)+M)^{k/h(k)}$
$\leq 2^{o(N)}+Nk^{k/h(k)}(2^{2N/k}+w+M)^{k/h(k)}$
$=2^{o(N)}+2^{o(N)}(2^{2N/k}+w+M)^{k/h(k)}\qquad\text{using }k\leq 2\log N$
$=2^{o(N)},$
where in the last step, we used the fact that if $w=2^{\lceil N/k\rceil}\geq M$ then $(2^{2N/k}+w+M)^{k/h(k)}\leq(3\cdot 2^{2N/k})^{k/h(k)}=2^{o(N)}$, otherwise $(2^{2N/k}+w+M)^{k/h(k)}=(O(N))^{k/h(k)}=2^{o(N)}$. ∎
5.3 NP-hardness of approximation for Uniform CapSoR
In this section, we prove Theorem 1.
Proof of Theorem 1.
We show the following reduction that is similar in construction to Lemma 13.
Lemma 20.
There is a reduction algorithm that for given an instance $\phi$ of Max3333-SAT5555 on $N$ variables and $M$ clauses produces an instance $\mathcal{I}=((P,\delta),k,U)$ of uniform CapSoR in time $N^{O(1)}$ such that the following holds.
  • $|P|=5N^{2}+11N/3$
  • $U=10N+7$
  • $k=N$
  • If $\phi$ is satisfiable then there exist a feasible solution $(S,\sigma)$ such that ${\rm cost}(S,\sigma)\leq k$
  • If every assignment satisfies less than $(1-\varepsilon^{\prime})M$ clauses, for some fixed $\varepsilon^{\prime}>0$, then for every feasible solution $(S,\sigma)$ it holds that ${\rm cost}(S,\sigma)<(1+\varepsilon^{\prime}/66)k$.
Proof Sketch.
We create an instance $\mathcal{I}=((P,\delta),F,k,U)$ of uniform CapSoR from $\phi$ by applying Lemma 13 using $k=N$. Note that, since $w=10N$, the capacity of every point $U=7+w=10N+7$ and the number of points in $\mathcal{I}$ is $|P|=N(2+w)+m=5N^{2}+2N+M=5N^{2}+11N/3$, as $M=5N/3$. The running time of the reduction is $N^{O(1)}$. The completeness and soundness follows from Lemma 14 and Lemma 15, respectively.
This finishes the proof of Theorem 1. ∎

### v2

@@ 6.2 Reduction from Max $k$-Cov to Uniform CapSoR
If $\exists\mathcal{T}=\{T_{1},\dots,T_{k}\mid T_{i}\in\mathcal{C}_{i},\forall i\in[k]\}$ such that $\mathcal{T}$ covers $\mathcal{G}$, then there exist a feasible solution $(S,\sigma)$ to the uniform CapSoR instance $\mathcal{J}$ with ${\rm cost}(S,\sigma)\leq k$.
@@
Without loss of generality, assume for every $T_{i}\in\mathcal{T},i\in[k]$, it holds that $T_{i}\in\mathcal{C}_{i}$. Consider the following solution $(S_{\mathcal{T}},\sigma)$ to $\mathcal{J}$, obtained from $\mathcal{T}$. Let $S_{\mathcal{T}}=\{p_{T_{1}},\dots,p_{T_{k}}\}$ contains all the set points corresponding to the sets in $\mathcal{T}$. For $p_{T}\in P$, where $p_{T}$ is a set point corresponding to set $T\in\mathcal{C}_{i},i\in[k]$, set $\sigma(p_{T})=p_{T_{i}}\in S_{\mathcal{T}}$ Note that this is feasible since $\mathcal{T}$ contains a set from each $\mathcal{C}_{i}$. Furthermore, since $\mathcal{T}$ covers $\mathcal{G}$, associate each point $e\in\mathcal{G}$ with some set $T_{j}\in\mathcal{T},j\in[k]$ covering $e$, by breaking ties arbitrarily. Call $T_{j}$ as the responsible set for $e$ in solution $\mathcal{T}$. Now, for an element point $p_{e}\in P$ corresponding to element $e\in\mathcal{G}$, set $\sigma(p_{e})=p_{T_{i}}\in S_{\mathcal{T}}$ such that $T_{i}$ is the responsible set for $e$ in $\mathcal{T}$. By construction of $(S_{\mathcal{T}},\sigma)$, the radius of each center in $p_{T_{i}}\in S_{\mathcal{T}}$ is $1$, and by the structure of $\mathcal{I}$, it holds that $\sigma^{-1}(p_{T_{i}})=\nicefrac{{n^{\prime}}}{{k}}+m=U$. Hence, $(S_{\mathcal{T}},\sigma)$ is a feasible solution to $\mathcal{J}$ with ${\rm cost}((S_{\mathcal{T}},\sigma))=k$. ∎
@@
If $\forall\mathcal{T}=\{T_{1},\dots,T_{k}\mid T_{i}\in\mathcal{C}_{i},\forall i\in[k]\}$, $\mathcal{T}$ covers strictly less than $(1-1/e+\varepsilon)n^{\prime}$ elements of $\mathcal{G}$, then then for every feasible solution $(S,\sigma)$ it holds that ${\rm cost}(S,\sigma)>(1+1/e-\varepsilon)k$.
@@
We prove the contraposition. Suppose there exists $S\subseteq P,|S|\leq k$ and a feasible assignment $\sigma:P\rightarrow S$ such that ${\rm cost}(S,\sigma)\leq(1+1/e-\varepsilon)k$. We will show that there exists $\mathcal{T}=\{T_{1},\dots,T_{k}\mid T_{i}\in\mathcal{C}_{i},\forall i\in[k]\}$ such that $\mathcal{T}$ covers at least $(1-1/e+\varepsilon)n^{\prime}$ elements of $\mathcal{G}$.
First note that every cluster in the solution $(S,\sigma)$ has precisely $n/k=U$ points, since $n=n^{\prime}+mk=Uk$ and each point has a capacity of serving at most $U=n^{\prime}/k+m$ points. Thus, we have that $|S|=k$ and there are no zero radius clusters in $(S,\sigma)$. Now, let $S_{1}\subseteq S$ be the centers that have radius $1$. We claim that $|S_{1}|\geq(1-1/e+\varepsilon)k$. Suppose $|S_{1}|<(1-1/e+\varepsilon)k$. Then, we have
${\rm cost}((S,\sigma))\geq|S_{1}|\cdot 1+(k-|S_{1}|)2>(1+\nicefrac{{1}}{{e}}-\varepsilon)k,$
which is a contraction to the cost of $(S,\sigma)$. Furthermore, since each cluster in the solution has $n^{\prime}/k+m$ points, we claim that $S_{1}$ contains at most one set point from each $\mathcal{C}_{j}$, and contains no element points. For the former, note that any two set points from $\mathcal{C}_{j},j\in[k]$ can cover at most $2n^{\prime}/k+m<2U$ points within distance $1$, and hence, one of them contains less than $U$ points. Similarly, for the latter, note that any element point can cover at most $mk+1\leq\nicefrac{{n^{\prime}}}{{2k}}+1<U=\nicefrac{{n^{\prime}}}{{k}}+m$ points within distance $1$, since $m\geq 1$. Now, consider the sets $\mathcal{T}=\{T_{1},\dots,T_{k}\}$ corresponding to the set points of $S_{1}$. Since, each set point of $S$ covers at most $m$ set points and in total precisely $U$ points, we have that $\mathcal{T}$ covers at least $|S_{1}|(U-m)\geq(1-\nicefrac{{1}}{{e}}+\varepsilon)n^{\prime}/k$ elements of $\mathcal{G}$, as desired.
This concludes the proof of the theorem. ∎
Finally, to obtain the proofs of Theorem 1 and 2, we use the reduction of Theorem 9 on the hard instances of Max $k$-Cov obtained from Theorem 7 and 8, respectively. Note that the condition $m<\nicefrac{{n^{\prime}}}{{2k^{2}}}$ in Theorem 9 can be easily achieved by duplicating $O(k^{2}m)$ times every element of $\mathcal{G}$. More specifically, in this case we duplicate each element $2mk^{2}$ times. Then, letting $n^{\prime\prime}$ denote the number of elements in the new instance, we have $n^{\prime\prime}=n^{\prime}\cdot 2mk^{2}$. Now, applying the reduction of Theorem 9 on this new instance of Max $k$-Cov, it implies that $U:=\nicefrac{{n^{\prime\prime}}}{{k}}+m$ is still a feasible constraint since each set now contains exactly ${2mk^{2}n^{\prime}}/{k}=n^{\prime\prime}/k$ elements. However, we also have that $m=\nicefrac{{n^{\prime\prime}}}{{2k^{2}n^{\prime}}}\leq\nicefrac{{n^{\prime\prime}}}{{2k^{2}}}$, as required. Finally, observe that since each element occurs $2n^{\prime}mk^{2}$ times, any solution to the new instance of Max $k$-Cov covering $\delta$ fraction of elements, also covers $\delta$ fraction of the elements in the original instance.
Acknowledgments.
We thank the anonymous reviewer for suggesting improved hardness of approximation construction.